RC Meeting, Wednesday, October 27 , 2021
RC Meeting Information
RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920
- Beam delivery back to Halls B,C,D on owl shift.
Hall C Status
Back in production at 42 muA on late swing. (Some beam early on swing, but mostly tuning for Hall B.)
On late swing, alarms due to dropping HB LHe levels. We have our good days, but they always come back. Steve L intervened. https://logbooks.jlab.org/entry/3933018 https://logbooks.jlab.org/entry/3933051 He's set it so the HB is now over-filling a bit.
End of swing, SHMS Q2 Interlock plus some sort of communication reset required on a screen I haven't seen before. https://logbooks.jlab.org/entry/3933073 Also Steve L. Recovery by 00:30 on owl. I didn't see this mentioned on our Wiki, so I added it, and listed it under Useful Links as well.
Earlier in the day, when there was no beam, Q3 slow dumped due to a communication failure. It took more than the usual power cycle to get it back: a small power supply had died and was replaced by Joe B and Steve L(?). https://logbooks.jlab.org/entry/3932750 https://logbooks.jlab.org/entry/3932747
Number of good pions is 5483.
Current Run Plan
- Day shift only: Try to take 1 hour of Dummy target each day. (a 100% live hour)
- BCM calibration: I would like to push this to Thursday, if the PD and other halls are willing. We'd really, really like 60+ muA so we can keep the errors on the fitted gains small, but I have no confidence the accelerator can do that today.
- Continue data taking until early Nov
- Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
- HMS Q3 insulating vacuum spikes generating annoyance alarms that take minutes to recover. https://logbooks.jlab.org/entry/3930175 Noise levels on transducers look OK, but Junaid's plot indicates Q3 is burping: https://logbooks.jlab.org/entry/3932992 .
- Serious HB LHe low liquid level alarms. https://logbooks.jlab.org/entry/3933051 Always seems to start with a subtle, simultaneous increase in level in Q1/2/3. https://logbooks.jlab.org/entry/3929152 .
- Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
- Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
- HCCAM02 not functioning (lift required to replace).
- How the harp scans and carbon hole target should look (see links):
- carbon hole target beam centering procedure:
- Resetting the "ioc heartbeat epics communications" (start by clicking on the heart at upper left of the magnet controls screen
- Fall 2021 Shift Schedule
- Hall-C Live Status
- Hall-C Logbook
- MCC White Board